OpenStudy (anonymous):

16*1+4(18/2-9) please show work.

OpenStudy (angela210793):

16+4*0=16+0=16

OpenStudy (stormfire1):

Just remember your order of operations and do the arithmetic...you shouldn't need help with this :)

OpenStudy (anonymous):

16+4(9-9) 16+4(0) 16+0 16